# Provable Ordering and Continuity in Vision-Language Pretraining for Generalizable Embodied Agents

### v1

@@ 4.2 Vision-Language Continuity
While the VLO property provides a strong global constraint on the structural alignment of VL pretraining, optimizing triplet relationships alone can be unstable. Variations in frame content and noise often lead to suboptimal local consistency. To mitigate this, we introduce an additional local continuity constraint inspired by the Brownian bridge (Revuz & Yor, 2013). This stochastic process models transitions between two fixed endpoints over by any sampled local video interval $[n(i),n(j)]$. For any time step $t\in[n(i),n(j)]$ within this interval, the transition density of Brownian Bridge process $\mathbf{B}(t)$ follows a time-dependent Gaussian distribution:
$\small\mathcal{N}\left(\mathbf{v}_{i}+\frac{t-n(i)}{n(j)-n(i)}(\mathbf{v}_{j}-\mathbf{v}_{i}),\frac{t(n(j)-n(i))-t^{2})}{n(j)-n(i)}\right),$
where $\mathbf{v}_{i},\mathbf{v}_{j}\in\mathbb{R}^{d}$ are the visual embeddings of the first and last frames in the sampled interval. The mean trajectory $\mathbb{E}[\mathbf{B}(t)]$ linearly interpolates between the two endpoints, while the variance $\mathrm{Var}[\mathbf{B}(t)]$ peaks provides uncertainty modeling that peaks in the middle of the interval. To enforce this local continuity, the Brownian bridge loss $\mathcal{L}_{\operatorname{BB}}$ is formulated as,
@@
Theoretical Analysis. To provide a deeper understanding of continuity preservation, we present theoretical guarantee:
Theorem 4.3 (Vision-Language Continuity).
Let $\mathbf{v}_{k},\mathbf{v}_{l}$ be arbitrary time step from the interval $[n(i),n(j)]$ and $\mathbf{l}\in\mathcal{L}$ be the language embedding. Suppose the VL similarity function $\operatorname{sim}(\cdot)$ is Lipschitz continuous with constant $C$. Assume that the frame embeddings are regularized by the Brownian Bridge constraint, then for any $\epsilon>0$, there exists $\delta>0$ such that
$\|\mathbf{v}_{k}-\mathbf{v}_{l}\|_{2}<\delta\Rightarrow\left|\mathfrak{R}(\mathbf{v}_{k},\mathbf{v}_{l},\mathbf{l})\right|<\epsilon.$
Proof.
Please refer to Appendix A.2. ∎
This theorem guarantees that under the Brownian Bridge constraint, if two video frames are temporally close, their semantic alignment score remains within a bounded range. This continuity can be furthur extended and show model resiliency to language perturbation:
Theorem 4.4 (Robustness to Language Variations).
Let $\mathbf{l}^{\prime}$ be the perturbed version of the original language embedding $\mathbf{l}$ subject to a small constant $\delta_{l}>0$, i.e., $\|\mathbf{l}-\mathbf{l}^{\prime}\|\leq\delta_{l}$, then the semantic alignment score $\mathfrak{R}$ exhibits stability to the perturbation:
$|\mathfrak{R}(\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{l}^{\prime})-\mathfrak{R}(\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{l})|\leq 2C\delta_{l}.$ (6)
Proof.
Please refer to Appendix A.3. ∎
Our empirical results, presented in Section 5.6, further validate these properties by showing improved generalization to varying linguistic instructions and more stable performance compared to baseline methods.
@@ 5 Experiment
Figure 2: Simulation environments, including two camera views of Franka Kitchen (5 tasks) and Metaworld (4 tasks).
Figure 3: The averaged success rate comparisons over various tasks on (a) Franka Kitchen and (b) Metaworld.
@@
Experimental Setup. We initialize our model with the weights of CLIP  (Radford et al., 2021) with ResNet50 vision backbone and further pre-train it on the large-scale human action video dataset EPIC-KITCHEN-100  (Damen et al., 2018, 2020). For hyperparameter selection, we randomly sample 10 frames from each video per batch. The loss weight $\lambda$ to 100. Other hyperparameters like temperatures follows the default value used in CLIP  (Radford et al., 2021). More pre-training details can be referred to Appendix B. In our experiments, we aim to evaluate the effectiveness of ordered and continuous vision-language representations for robotic control. First, we conduct extensive Language-Conditioned Behavior Cloning (LCBC) experiments in two different simulation environments to validate the importance of ordering and continuity for imitation learning. Second, we assess the utility of the learned representations as reward functions on multiple real-world action videos. The results demonstrate that the ordered and continuous representations enable our method to accurately identify action boundaries and generate dense rewards aligned with the given instructions. Finally, we evaluate the robustness of our method under language perturbations, showcasing its strong generalization capability for application in real-world daily scenarios.
Figure 4: Visualization of the normalized learned reward corresponding to different actions. Our representations effectively help capture the correct temporal order of actions in the instruction. For more results, please refer to Appendix C.4.
5.1 Simulation Environments
We perform LCBC experiments in two widely used simulation environments for evaluation: Franka Kitchen  (Gupta et al., 2019; Fu et al., 2020) and Metaworld  (Yu et al., 2019). As shown in Figure  2, for Franka Kitchen, we evaluate five tasks: sliding a cabinet, opening the left door, opening the microwave, turning on the stove, and switching on the light. For Metaworld, we focus on learning four tasks: hammering a nail, pressing a button, picking and placing a block, and assembling a ring onto a peg. Detailed environment setup can be found at Appendix C.1.
In Frankakitchen, tasks often involve intricate actions that induce large visual changes, and the successful completion of these tasks requires precise, complex actions. As a result, tasks on FrankaKitchen rely more on visual trajectory representations. On the other hand, the visual scene in Metaworld is much simpler. Tasks on Metaworld generally involve direct actions with smaller visual changes, yet they require more sophisticated language understanding to act towards specific task goals.
5.2 Baselines
Since our model is initialized with CLIP  (Radford et al., 2021), a state-of-the-art image-text representation widely applied in various embodied tasks  (Cui et al., 2022; Khandelwal et al., 2022; Shridhar et al., 2021; Tam et al., 2022), it is a natural choice to include CLIP as a vanilla baseline for comparison. Our primary baselines are LIV  (Ma et al., 2023a) and DecisionNCE  (Li et al., 2024), as we all use the same model architecture and dataset for pretraining. LIV employs the VIP  (Ma et al., 2023b) to achieve consistent frame representations and aligns the final frame with instructions using the CLIP loss. DecisionNCE represents instructions as frame transitions and aligns the difference between the initial and final frames with the instructions using the CLIP loss. We also compare against R3M  (Nair et al., 2022), pre-training on Ego4D  (Grauman et al., 2022), which combines time contrastive learning  (Sermanet et al., 2018) with LOReL  (Nair et al., 2021) to ensure that later frames in the sequence receive higher rewards when aligned with the instruction.
@@ 5.3 Language-Conditioned Behavior Cloning
We keep the pre-trained vision-language encoders frozen and feed their output representations into a lightweight MLP to train the LCBC policies. Each task is performed from two camera viewpoints (left and right), with varying numbers of demonstrations $\left[5,15,25\right]$ (i.e., dataset size) for training, and evaluated under three different random seeds. We report the success rate across different environments and dataset sizes, averaged over camera views and seeds. Detailed comparison results can be referred to Appendix C.3.
Figure 3 presents the comparison results, showing that AcTOL significantly outperforms other methods across all dataset sizes in both environments, surpassing the SOTA by 49.0%/46.4%/26.3% and 14.6%/15.2%/16.3% at dataset sizes of 5/15/25, respectively, which demonstrates the effectiveness of our pre-training strategy. Especially in Franka Kitchen, where complex tasks demand higher action continuity, our method demonstrates a clear advantage with limited data, highlighting its superior data efficiency and low-resource generalization capability. Among the baseline models, CLIP consistently shows lower performance, particularly for Franka Kitchen. This indicates that while it excels in image representation, its inability to capture temporal characteristics in videos leads to suboptimal results in control tasks. R3M, LIV, and DecisionNCE perform well in the FrankaKitchen environment, but R3M shows a significant drop in performance in Metaworld. This discrepancy stems from R3M’s design, which utilizes a frozen language encoder to generate language features as conditions. These features are then used to optimize the visual encoder by maximizing the reward for temporally later frames. This approach results in a shortcut that prioritizes maximizing temporal visual differences rather than genuinely aligning visual representations with language semantics. This limitation is particularly evident in Franka Kitchen, where large visual changes dominate, allowing R3M to excel despite its lack of true language-visual alignment. In contrast, LIV and DecisionNCE employ CLIP-based losses during pre-training on video data, enabling them to align visual changes with semantics. This capability allows them to maintain solid performance even in Metaworld, where tasks require more refined language comprehension to interpret subtle visual changes. However, both models suffer from rigid assumptions about action semantics during pre-training, resulting in suboptimal alignment between vision and instructions.
Notably, when using 5/15 demonstrations, our method achieves results that are comparable to, or even surpass, those of other methods using 15/25 demonstrations, illustrating that our approach can utilize expert data more effectively. This is particularly advantageous when collecting expert data is time-consuming and labor-intensive. Moreover, when the Brownian Bridge constraint is not applied, although our method still outperforms the other baselines, its performance decreases noticeably. This indicates that the Brownian Bridge constraint effectively improves the quality of learned representations, leading to better policy optimization in behavior cloning.
@@ 5.4 Language-Conditioned Visual Rewards
Since our model learns semantically smooth visual representations, the resulting semantic trajectories can also serve as ideal task rewards. Specifically, we define the reward at time step $i$ as $\operatorname{cosine}(\mathbf{v}^{i},\mathbf{l})$, which reflects the distance between the current state and the language goal. Previous works  (Ma et al., 2023a; Li et al., 2024) have primarily tested their rewards on single-action video clips. To increase task complexity, we selected three video clips, each containing two consecutive actions, to better evaluate whether the model accurately understands action semantics. We present more reward results in Appendix C.4. Figure  4 shows the actions, the instructions given and the corresponding reward curves. Figure  4(a) uses a video from EPIC-KITCHEN-100 to evaluate the effectiveness of in-distribution rewards. Our method produces an ideal reward curve, starting at a low point and reaching a peak as the action “open cupboard” completes, before declining as subsequent actions begin. In contrast, we observe that the rewards generated by R3M and DecisionNCE continue to rise even after the “open cupboard” action has finished. This indicates that their training methodologies fail to effectively distinguish between actions that align with the instruction and those that do not. Figure  4(b) and (c) evaluate the reward generation capabilities in real-world scenarios using two videos from  (Bahl et al., 2022). These videos involve humans and robots consecutively performing pairs of opposite actions. It is evident that only our method is capable of generating the correct reward curve, accurately identifying the start and end of the actions specified by the instruction, and producing opposite rewards for contrasting actions. This demonstrates that our method, by learning ordered and continuous vision-language representations, inherently captures the trends of semantic changes in actions. It effectively aligns visual trajectories with the corresponding instructions, highlighting its significant potential as a language-conditioned visual reward model.
Figure 5: Visualization of visual trajectory representations.
5.5 Visualization of Visual Representation Trajectory
To demonstrate the smoothness of the representations, we select three distinct language instructions from the EPIC-KITCHEN-100 dataset, each corresponding to 10 action videos. We then visualize the visual representations learned by our method and those by CLIP using t-SNE  (van der Maaten & Hinton, 2008), as shown in Figure  5. The representations produced by CLIP show clear separability between different language instructions, and the trajectories of individual videos maintain a certain degree of temporal consistency. However, the transitions between consecutive frames are not smooth, reflecting a flaw due to the absence of training on video data. In contrast, our method significantly enhances the ordering and continuity of video feature trajectories while preserving the discriminative power of CLIP for distinguishing actions associated with different instructions. This improvement stems from training directly on CLIP’s weights and optimizing the temporal consistency within each video. As a result, our method not only achieves smoother representations but also retains the strong alignment between visual features and language semantics inherent in CLIP’s original design. This balance between temporal coherence and semantic distinctiveness highlights the effectiveness of our approach in refining embodied representations for tasks requiring nuanced understanding of multimodal alignment and temporal dynamics.

### v3

@@ 3.2 Vision-Language Continuity
While the VLO property provides a strong global constraint on the structural alignment of VL pretraining, optimizing triplet relationships alone can be unstable. Variations in frame content and noise often lead to suboptimal local consistency. To mitigate this, we introduce an additional local continuity constraint inspired by the Brownian bridge [revuz2013continuous]. This stochastic process models transitions between two fixed endpoints over by any sampled local video interval $[n(i),n(j)]$. For any time step $t\in[n(i),n(j)]$ within this interval, the transition density of Brownian Bridge process $\mathbf{B}(t)$ follows a time-dependent Gaussian distribution:
$\mathcal{N}\left(\mathbf{v}_{i}+\frac{t-n(i)}{n(j)-n(i)}(\mathbf{v}_{j}-\mathbf{v}_{i}),\frac{t(n(j)-n(i))-t^{2})}{n(j)-n(i)}\right),$
where $\mathbf{v}_{i},\mathbf{v}_{j}\in\mathbb{R}^{d}$ are the visual embeddings of the first and last frames in the sampled interval. The mean trajectory $\mathbb{E}[\mathbf{B}(t)]$ linearly interpolates between the two endpoints, while the variance $\mathrm{Var}[\mathbf{B}(t)]$ provides uncertainty modeling that peaks in the middle of the interval. To enforce this local continuity, the Brownian bridge loss $\mathcal{L}_{\operatorname{BB}}$ is formulated as,
@@
4 Theoretical Analysis
In this section, we theoretically prove the vision-language ordering and continuity, as well as extend the robustness of linguistic perturbations of representations learned by AcTOL. All proofs are provided in Appendix C for reference.
Vision-Language Ordering. Ordering and sorting properties are well-established in self-supervised learning [DBLP:conf/iccv/ShvetsovaPKSK23, DBLP:conf/iccv/HuSLRSS21, nips23-rnc]. Building upon these insights, we formalize the concept of vision-language ordering (VLO) below.
Definition 1 (VLO Representations).
[…]
To formalize the temporal ordering constraints, we define the unique sorted set of frame distances from frame $i$ as $\{D_{i,1}<D_{i,2}<\cdots<D_{i,M_{i}}\}$, where each $D_{i,m},m\in[M_{i}]$ is obtained by sorting the set $\{d_{i,j}\mid j\in[T]\setminus\{i\}\}$. Additionally, we define the count of frames at each distance level as:
$n_{i,m}:=|\{j\mid d_{i,j}=D_{i,m},\,j\in[T]\setminus\{i\}\}|,$ (4)
This bound characterizes the optimal alignment of VL similarities, ensuring that the learned representations preserve the inherent temporal structure within the video sequence, as guaranteed by the following theorem:
Theorem 1 (Vision-Language Ordering).
$\mathcal{L}^{*}$ is a tight lower bound of $\mathcal{L}_{\mathrm{VLO}}$, i.e., $\mathcal{L}_{\mathrm{VLO}}\geq\mathcal{L}^{*}$, and for any $\epsilon>0$, there exists feature embeddings such that $\mathcal{L}_{\mathrm{VLO}}<\mathcal{L}^{*}+\epsilon$. Furthermore, for any $0<\delta<1$, there exist $\epsilon>0$ such that if $\mathcal{L}_{\mathrm{VLO}}<\mathcal{L}^{*}+\epsilon$, the learned representations satisfy the VLO property.
Vision-Language Continuity. We establish the following theoretical result to rigorously describe continuity preservation in vision-language representations:
Theorem 2 (Vision-Language Continuity).
Let $\mathbf{v}_{k},\mathbf{v}_{l}$ be visual representations at arbitrary time steps within a Brownian Bridge-regularized interval $[n(i),n(j)]$, and let $\mathbf{l}\in\mathcal{L}$ be a language embedding. If the VL similarity function $\operatorname{sim}(\cdot)$ is Lipschitz continuous with constant $C$, then for any $\epsilon>0$, there exists $\delta>0$ such that: $\|\mathbf{v}_{k}-\mathbf{v}_{l}\|_{2}<\delta\quad\Rightarrow\quad\left|\mathfrak{R}(\mathbf{v}_{k},\mathbf{v}_{l},\mathbf{l})\right|<\epsilon.$
This result follows from two key observations: (i) Brownian Bridge regularization constrains each embedding to remain close to a linear interpolation between anchor frames, with deviations governed by a time-dependent variance; and (ii) under this constraint, the distance between temporally close frames admits an explicit upper bound. Combining this with the Lipschitz continuity of the vision-language similarity function ensures that small changes in frame embeddings lead to proportionally bounded changes in alignment scores.
Building upon the continuity result, we further demonstrate that the semantic alignment score remains stable under small perturbations in language input:
Theorem 3 (Robustness to Language Variations).
Let $\mathbf{l}^{\prime}$ be a perturbed language embedding such that $\|\mathbf{l}-\mathbf{l}^{\prime}\|\leq\delta_{l}$. Then the semantic alignment score $\mathfrak{R}$ satisfies: $|\mathfrak{R}(\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{l}^{\prime})-\mathfrak{R}(\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{l})|\leq 2C\delta_{l}.$
This second result guarantees that small shifts in the language representation (e.g., synonym substitution or phrasing variation) lead to bounded changes in the alignment score. Together, Theorems 2 and 3 formalize the local stability of semantic grounding across both time and modality, providing a theoretical basis for continuity-aware vision-language learning.
@@ 5 Experiment
In our experiments, we aim to evaluate the effectiveness of ordered and continuous vision-language representations for robotic control. First, we conduct extensive Language-Conditioned Behavior Cloning (LCBC) experiments on both real and simulated robots to validate the importance of ordering and continuity for imitation learning. Second, we assess the utility of the learned representations as reward functions on multiple real-world action videos. The results demonstrate that the ordered and continuous representations enable our method to accurately identify action boundaries and generate dense rewards aligned with the given instructions. Finally, we evaluate the robustness of our method under language perturbations, showcasing its strong generalization capability for application in real-world daily scenarios.
Figure 3: Policy learning environments, including 3 tasks with a real-world Unitree D1 robot arm and 5 tasks each in two simulation environments, i.e., Franka Kitchen and Metaworld.
Experimental Setups. Figure 3 shows the experimental environments. For real-world robot evaluation, we deploy the Unitree D1 robot arm to perform three challenging manipulation tasks: pick cup, open [X] drawer and close [X] drawer, where [X] is the drawer index specified by the instruction. The pick cup task requires the model to accurately identify the cup handle, while the open/close [X] drawer tasks demand grounding of language instructions to visual observations, enabling the model to interact with the correct drawer. To isolate manipulation performance, the Unitree Go2 quadruped remains lying down and stationary throughout the evaluation. We use a web camera to capture a third-person view as visual observation. The action space consists of a 6-DoF end-effector displacement vector and gripper state, executed at a control frequency of 20 Hz. For each task, we collect 60 demonstrations via remote control using the Unitree Go app, which is significantly fewer than the 100 trajectories typically used in prior work [icml23-liv, icml24-decisionnce]. For simulation, we choose two widely used simulation environments for evaluation: Franka Kitchen [corl19-franka, corr20-franka] and Metaworld [corl19-metaworld]. For Franka Kitchen, we evaluate five tasks: sliding a cabinet, opening the left door, opening the microwave, turning on the stove, and switching on the light. For Metaworld, we focus on learning five tasks: hammering a nail, pressing a button, picking and placing a block, assembling a ring onto a peg, and opening a drawer. Detailed environment setup can be found at Appendix B.1.
Baselines. Since our model is initialized with CLIP [icml21-clip], a state-of-the-art image-text representation widely applied in various embodied tasks [l4dc22-clipapp2, cvpr22-clipapp3, corl21-clipapp1, nips22-clipapp4], it is a natural choice to include CLIP as a vanilla baseline for comparison. Our primary baselines are LIV [icml23-liv] and DecisionNCE [icml24-decisionnce], as we use the same model architecture and dataset for pre-training. We also compare against R3M [corl22-r3m] pre-trained on Ego4D [cvpr22-ego4d], a dataset containing roughly $36\times$ longer videos than EPIC-KITCHEN-100. We also include an ablation variant of AcTOL where the Brownian Bridge loss is removed, referred to as AcTOL w/o BB.
Implementation Details. We initialize our model with the weights of CLIP [icml21-clip] with ResNet-50 vision backbone and further pre-train it on human action video dataset EPIC-KITCHEN-100 [corr18-epickitchen, corr20-epickitchen]. For hyperparameter selection, we uniformly sample 10 frames of each video per batch. The loss weight $\lambda$ is 0.1. Other hyperparameters, such as temperature,s follow the default value used in CLIP [icml21-clip]. More details of pre-training and hyperparameter sensitivity can be found in Appendix A.
Table 1: Comparison in simulation environments with varying amounts of demonstrations. Each result reports the success rate over 50 roll-outs, averaged across 2 camera views and 3 random seeds. We also report the relative performance gain in green compared to the strongest baseline.
Method Franka Kitchen Metaworld
5 demos 15 demos 25 demos 5 demos 15 demos 25 demos
CLIP 11.67 ± 0.95 27.47 ± 1.01 31.20 ± 2.62 42.29 ± 2.65 60.33 ± 1.32 62.54 ± 4.36
R3M 28.60 ± 1.39 42.20 ± 1.00 51.13 ± 2.83 46.83 ± 3.85 56.50 ± 5.20 60.08 ± 3.62
LIV 23.40 ± 0.78 42.73 ± 1.17 51.93 ± 0.95 46.95 ± 2.07 64.33 ± 3.63 66.67 ± 1.49
DecisionNCE 25.33 ± 1.30 43.20 ± 2.25 50.87 ± 2.95 44.58 ± 2.79 59.08 ± 1.77 69.75 ± 3.90
AcTOL w/o BB 32.80 ± 1.23 54.20 ± 0.85 60.80 ± 0.87 50.29 ± 4.05 70.83 ± 4.21 73.33 ± 2.83
AcTOL 42.60 ± 0.53 61.80 ± 2.54 64.60 ± 0.57 53.81 ± 3.89 74.13 ± 1.59 81.13 ± 1.59
(+48.95%) (+43.06%) (+24.40%) (+14.61%) (+15.23%) (+16.32%)
@@ 5.1 Language-Conditioned Behavior Cloning
For LCBC policy learning, we keep the pre-trained vision-language encoders frozen and feed their output representations into a lightweight MLP, which is trained as a policy network.
Simulation results. In simulation, each task is performed from two camera viewpoints (left and right), with varying numbers of demonstrations $\left[5,15,25\right]$ (i.e., dataset size) for training, and evaluated under three different random seeds. We report the success rate across different environments and dataset sizes, averaged over camera views and seeds. Detailed comparison results for each task can be referred to Appendix B.5. Table 1 presents the comparison results, demonstrating that AcTOL achieves significantly enhanced performance relative to baseline methods across all evaluated datasets and environments. This superiority is particularly pronounced in the complex Franka Kitchen setting, especially under data constraints, where AcTOL with fewer demonstrations (e.g., 5/15) often matches or surpasses other methods using more data (e.g., 15/25), indicating its high data efficiency and robust low-resource generalization capabilities. Furthermore, ablation studies confirm the integral role of the Brownian Bridge (BB) constraint; its removal (AcTOL w/o BB) results in a significant performance decrease, validating its contribution to improving representation quality for effective policy optimization via behavior cloning.
Figure 4: Visual shifts applied in Franka Kitchen.
Robustness under visual shifts. To further assess the ability of the model to handle visual distribution shifts, we conduct experiments in the Franka Kitchen environment following the protocol of [burns2023makespretrainedvisualrepresentations]. Specifically, we compare AcTOL with the strongest baseline, DecisionNCE, under visual changes absent from training. These include: (1) object distractors of increasing difficulty: easy (S1), medium (S2), and hard (S3), corresponding to scenes with 1, 3, and 9 additional YCB objects [calli2015ycb], respectively; and (2) background texture variations with marble hinge (S4) and metal slide (S5). All shifts are shown in Figure 4.
Table 2: Success rate comparison across different visual shifts in Franka Kitchen.
Method No shift S1 S2 S3 S4 S5
DecisionNCE 43.2 27.2 25.6 4.8 0.0 8.8
AcTOL 61.8 43.2 32.8 9.2 4.4 38.4
Policies are trained with 15 demonstrations per task, and success rates averaged over five tasks are reported in Table 2. While performance drops under visual shifts, which is expected, AcTOL continues to outperform DecisionNCE in all available test conditions. This suggests that the learned representation maintains useful generalization ability even without any specific adaptation for visual domain shift.
Table 3: Performance comparison on Unitree D1 arm. Success rates are reported over 10 trials.
Method Pick Cup Open [X] Drawer Close [X] Drawer
CLIP 0 20 30
R3M 10 40 40
LIV 20 30 50
DecisionNCE 20 40 60
AcTOL 50 80 90
Real Robot results. Table 3 shows the real robot comparison results. AcTOL consistently outperforms all baseline models across the three tasks. Among them, the pick cup task yields relatively lower performance, as it requires the model to precisely identify and grasp the cup handle, demanding stronger spatial perception capabilities. For the open/close [X] drawer tasks, AcTOL is able to accurately interpret the drawer number specified in the language instruction, align it with the corresponding location in the visual observation, and execute continuous actions on the correct drawer to complete the task. These results highlight the effectiveness of AcTOL’s learned visual-language representations in real-world manipulation tasks.
Figure 5: Visualization of the normalized learned reward corresponding to different actions. Our representations effectively help capture the correct temporal order of actions in the instruction. For more results, please refer to Appendix B.6.
@@ 5.2 Language-Conditioned Visual Rewards
Figure 6: Visual trajectory visualization.
By learning semantically smooth visual representations, our model further enables the use of semantic trajectories as effective task rewards. To illustrate this, we first demonstrate the continuity of purely visual representations. In Figure 6, we visualize the learned visual representation trajectories for three tasks, each with ten video clips, using t-SNE. The results show that AcTOL significantly improves the temporal continuity of video feature trajectories while retaining CLIP’s discriminative ability to distinguish between actions associated with different instructions. As discussed in Section 3.2, the visual continuity can stabilize learning ordered vision-language alignment. Building on this foundation, we define a dense reward signal based on the semantic alignment between the current visual state and the language goal. Specifically, at each time step $i$, we define the reward $\operatorname{cosine}(\mathbf{v}^{i},\mathbf{l})$ as the similarity between the current visual state and the language goal. While prior work [icml23-liv, icml24-decisionnce] focused primarily on single-action video clips, we evaluate reward quality on three clips, each containing two consecutive actions, to assess whether the model can reliably capture fine-grained action semantics. Figure 5 (a) presents an in-distribution evaluation using a video from EPIC-KITCHEN-100. Our model produces a clear reward peak aligned with the completion of the “open cupboard” action, followed by a decline, indicating successful temporal localization of the instructed behavior. In contrast, R3M and DecisionNCE rewards continue increasing beyond the relevant action segment. Figures 5 (b) and (c) show results on real-world videos from [rss22-robotvideo], where human and robot actors perform opposite actions. Only our method consistently produces symmetric and instruction-aligned reward curves, accurately identifying both action boundaries and semantics.